# Occlusion-Aware Ground Target Tracking by a Dubins Vehicle Using Visibility Volumes

### v1

@@ 5.1 Feasibility of Reference Orbits
which is the value of $\kappa$ obtained by substituting (14) into (32), using the identities $\bm{g}\cdot\bm{e}_{r}=v_{g}\cos a$ and $\bm{g}\cdot\bm{e}_{\theta}=v_{g}\sin a$, defining $a=\theta-\gamma$ as the difference between the polar angle and the POI’s direction, introducing the speed ratio $b=\frac{v_{g}}{v}$, defining the ratio between the orbits radial rate and the UAV’s speed $c=\frac{\dot{R}}{v}$, the initial radius $d=R_{0}$, and $R=d+vct$ as the orbit radius at any time $t$. The bounds of the maximization are the bounds on the rate of change of the orbit’s radius from (53),
11 11 $\geq b+c$
$-1$ $\leq-b+c\;.$
The Jacobian of (55) was computed symbolically as
$\nabla J=\begin{bmatrix}\left(b\sin(a)\pm\sqrt{1-(b\cos(a)+c)^{2}}\right)\\ \Bigg{\{}\frac{-2b\left(\cos(a)\pm\frac{\sin(a)(b\cos(a)+c)}{\sqrt{1-(b\cos(a)+c)^{2}}}\right)}{(cvt+d)\left(1-(b\cos(a)+c)^{2}\right)^{1/2}}\\ +\frac{b\sin(a)(b\cos(a)+c)\left(b\sin(a)\pm\sqrt{1-(b\cos(a)+c)^{2}}\right)}{(cvt+d)\left(1-(b\cos(a)+c)^{2}\right)^{3/2}}\Bigg{\}}\\ \\ \left(b\sin(a)\pm\sqrt{1-(b\cos(a)+c)^{2}}\right)\\ \Bigg{\{}\frac{\pm 2(b\cos(a)+c)}{(cvt+d)\left(1-(b\cos(a)+c)^{2}\right)}\\ -\frac{(b\cos(a)+c)\left(b\sin(a)\pm\sqrt{1-(b\cos(a)+c)^{2}}\right)}{(cvt+d)\left(1-(b\cos(a)+c)^{2}\right)^{1/2}}\\ \pm\frac{vt\left(b\sin(a)\pm\sqrt{1-(b\cos(a)+c)^{2}}\right)}{(cvt+d)^{2}\left(1-(b\cos(a)+c)^{2}\right)^{1/2}}\Bigg{\}}\end{bmatrix}\;.$ (56)
Figure 3 illustrates the cost function (55) for counter-clockwise motion with parameters $d=1$, $t=0$, and a range of $b$ values. The curves where elements of the Jacobian (56) are zero also overlayed in Fig. 3.
Figure 3: A plot of the cost function (55) and its Jacobian (56) where $d=1$, $v=1$ and $t=0$ and for varying values of $b$. The value of the cost function is shown by the dark red to yellow map, and the locations where the elements of the Jacobian are zero are displayed as blue and green lines. The cyan dots are the locations where both elements of the Jacobian are zero, indicating extrema. The gray areas are invalid regions according to Lemma 5.
In Fig. 3, two local extrema are observed to always occur at $a=\pm\frac{\pi}{2}$ and $c=0$. For any set of $d,t$, there is an equivalent value of $d^{\prime},t^{\prime}=0$ such that $cvt+d=cvt^{\prime}+d^{\prime}$; Therefore, evaluating (56) at $t^{\prime}=0$ and $d^{\prime}$ is valid for the corresponding $t>0$ and $d$. Evaluating (56) at $a=\pm\frac{\pi}{2}$ and $c=0$ symbolically results in the zero vector and shows that the critical point is independent of $b$, $d$, and $v$. Thus, the function (55) has local extrema at [±$[\pm\frac{\pi}{2},0]^{\rm T}$ with a value of,
@@
Based on an analysis of plots, the maximum for counter-clockwise motion is at $\frac{\pi}{2}$ and $-\frac{\pi}{2}$ for clockwise motion. ∎
@@
Since $v_{g}\leq v$, then from Lemma 6, the maximum of the right-hand side of (54) is $\kappa=\frac{4}{R}$. This indicates that a UAV with speed $v$ can track a moving POI with any speed $v_{g}\leq v$ as long as the orbit radius is four times greater than the minimum radius $R\geq 4r_{\rm min}$. This is useful because it creates a UAV turn radius selection criterion for a tracking mission given the mission environment.
@@ 5.2 Representing Moving View Volumes via an Adaptive Discretization Approach
VVs for a moving POI remain relatively constant in shape over a road network when the surrounding building geometry also remains relatively unchanged (e.g., when traversing between two spatially uniform structures). However, when one structure ends or changes shape, the VV geometry varies. An adaptive discretization approach chooses an appropriate spacing over the road network for which to compute VVs such that more VVs are located in regions of greater environmental variability. The adaptive discretization approach relies on quantifying the difference between neighboring VVs $\hat{\alpha}(\bm{g}_{i})$ and $\hat{\alpha}(\bm{g}_{j})$, computed at two points $\bm{g}_{i}$ and $\bm{g}_{j}$, using the distance metric:
@@
The proposed adaptive approach is inspired by a bisection method and is described in Alg. 1. The algorithm increases the number of VVs along a road network edge where the VVs are quickly changing by placing a new VV halfway between two VVs. The inputs into the algorithm include the start and end points of an edge on the road network, $\bm{g}_{i}$ and $\bm{g}_{j}$, a set of initial points along the road network $\mathcal{P}$, and a constant threshold parameter $d_{\rm cutoff}$ that is repeatedly compared to (58). The output of the algorithm is a set of points $\mathcal{P}$ at which the VVs were computed. Initially, the algorithm initializes this set of points with $\mathcal{P}=\{{\bm{g}}_{i},{\bm{g}}_{j}\}$. Then, the recursive algorithm checks if the two input points along the edge produce two VVs where the difference metric is less than $d_{\rm cutoff}$. If that condition is true then the set of points $\mathcal{P}$ is unaltered and returned (Alg. 1, Lines 4-6). However, if the difference between VVs is larger than $d_{\rm cutoff}$, the point in the middle of the two input points $\bm{g}_{m}$ is created and added to the set of order points (Alg. 1 Lines 7-8). Then, the the algorithm proceeds in a recursive manner—first using $\bm{g}_{i}$ and $\bm{g}_{m}$ as the inputs (Alg. 1, Line 9), and then $\bm{g}_{m}$ and $\bm{g}_{j}$ as the inputs (Alg. 1, Line 10). The algorithm continues to subdivide each line segment until the threshold $d_{\rm cutoff}$ is met between each pair of consecutive points in $\mathcal{P}$. Finally, line 11 in Alg. 1 returns the updated set of ordered points.
Algorithm 1 Adaptive Discretization
1: function: AdaptiveDiscretization$(\bm{g}_{i},\bm{g}_{j},\mathcal{P},d_{\rm cutoff})$
2: input: $\bm{g}_{i}$ a point up the road network edge, $\bm{g}_{j}$ a point down the road network edge, $\mathcal{P}$ an ordered set of points along the road network edge, $d_{\rm cutoff}$ metric cutoff for two VVs
3: output: an ordered set of points along a road network edge $\mathcal{P}$
4:if $\bm{g}_{i}=\bm{g}_{j}$ or $d_{\rm cutoff}>d(\hat{\alpha}(\bm{g}_{i}),\hat{\alpha}(\bm{g}_{j}))$ then
5:   return $\mathcal{P}$
6:end if
7:$\bm{g}_{m}\leftarrow 0.5(\bm{g}_{i}+\bm{g}_{j})$
8:$\mathcal{P}\leftarrow\mathcal{P}\cup\bm{g}_{m}$
9:$\mathcal{P}\leftarrow\texttt{AdaptiveDiscretization}(\bm{g}_{i},\bm{g}_{m},\mathcal{P},d_{\rm cutoff})$
10:$\mathcal{P}\leftarrow\texttt{AdaptiveDiscretization}(\bm{g}_{m},\bm{g}_{j},\mathcal{P},d_{\rm cutoff})$
11:return $\mathcal{P}$
Figure 4: A plot of the VV similarity metric along a path through an urban environment for fixed distance discretization methods ($50~{}{\rm m}$ blue, $25~{}{\rm m}$ orange, $10~{}{\rm m}$ green, and $1~{}{\rm m}$ red) and the adaptive discretization approach (purple) with a cutoff at $d_{\rm cutoff}=500,000$ m3. This choice of cutoff is about 25% of the volume of a hemispherical VV with a radius of 100 m.
Figure 4 plots the distance (58) computed for a set of points $\mathcal{P}$ along a path through a city to compare the adaptive discretization with several fixed-spacing discretizations with uniform spacings of $1~{}{\rm m}$, $10~{}{\rm m}$, $25~{}\rm{m}$, $50~{}{\rm m}$. The metric (58) varies greatly for the larger spacings, but the adaptive sampling method maintains the metric at an approximately constant level or below $d_{\rm cutoff}=500,000$ m3. The value for the adaptive sampling is similar to the $10~{}{\rm m}$ spacing but reduces the spike around $1300~{}{\rm m}$ along the path. A uniform sampling with $1~{}{\rm m}$ spacing produces an even more refined representation of the VVs, but it requires more computation and memory storage.
@@ 5.3 Locus of Circular Orbits
Once the path of the POI on the road network has been discretized into a series of $m$ points $\mathcal{P}=\{{\bm{g}}_{0},\ldots,{\bm{g}}_{m-1}\}$ with corresponding VVs $\hat{\alpha}({\bm{g}}_{i})$ for all ${\bm{g}}_{i}\in\mathcal{P}$ at corresponding time instants $t_{0:m-1}=\{t_{0},\ldots,t_{m-1}\}$, then the locus of orbits is created with the set of radii $R_{0:m-1}=\{R_{0},\ldots,R_{m-1}\}$ with centers corresponding to points in $\mathcal{P}$. These radii define the piecewise-linear time-varying radius $R$ as described in Sec. 5.1. The objective (7) is to maximize the standoff distance to the target. Therefore, the radii $R_{i}$ are chosen to be as large as possible while remaining within the VPs (at altitude $h_{\rm UAV}$) corresponding to $\hat{\alpha}({\bm{g}}_{i})$ while being feasible for the Dubins vehicle with maximum curvature $\kappa_{\rm max}$.
@@
2: input: $\bm{g}(t)$ the POI trajectory, a set $\mathcal{P}$ of $m+1$ points along the trajectory $\bm{g}(t)$, $t_{0:m}$ the set of times corresponding to the points, $\kappa_{\rm max}$ the maximum curvature of the UAV, $v$ the UAV’s speed
3: output: an ordered set of radii $R_{0:m}$ where all points on the orbit can view the corresponding point in $\mathcal{P}$
@@
The algorithm first uses the bisection method MaxRadius to find the largest circle centered on $\bm{g}_{i}$ that can fit inside each VV $\hat{\alpha}({\bm{g}}_{i})$ at height $h_{\rm UAV}$ (Alg. 2, Line 4). The process of finding the largest VO is similar to the maximum inscribed circle problem [57, 58, 59] except that the center of the VO is fixed. The bisection method starts with the minimum radius 00 and the maximum sensing distance $d_{\rm max}$. The bisection method then checks to see if the radius halfway in between the minimum and maximum is inside the VP. If it is, the minimum radius is updated to the new intermediate radius; if not, the maximum radius is updated to the intermediate radius. The bisection method ends when the minimum and maximum radii are arbitrarily close, resulting in a circle that touches the VP in one or more places. Then, to ensure that the UAV can fly along the morphing orbit, the rate of change of the orbits is checked according to Lemma 5. The first for loop (Alg. 2, Lines 5-10) iterates forward through the orbits and decreases the next orbit if the radius increases too quickly. If there is a small orbit and then a series of large orbits, then the forward iteration will start decreasing the series of large orbits, obeying the upper bound from (51). If, during the forward loop, the lower bound of (51) is used to decrease the current radius, then the orbit between the preceding radius and the current radius may shrink too quickly. So, the second for loop (Alg. 2, Lines 11-16) iterates through the orbits backward and decreases the preceding orbit if the radius decreases too quickly. The backward iteration assures that the rate of change of the radii of preceding orbits is respected. Then, the resulting morphing orbit is checked for curvature constant violations (Alg. 2, Line 18) with Lemma 6. If the check for the curvature of the orbit fails, the path $\bm{g}$ is deemed infeasible for system (1)-(2) with speed $v$ and curvature bound $\kappa_{\rm max}$. An example of the resulting circular orbits can be seen in Fig. 5. Where the orbit is morphing from the $i$th orbit inside the $i$th VV to the $j$th orbit inside the $j$th VV. The $i$th and $j$th VVs are calculated along the POI’s path. Figure 5 also illustrates that a choice of circular orbits restricts the UAV to a small percentage of the visibility polygon for some environments. Future work may consider more general visibility orbit geometries.
Figure 5: A circular orbit inside the (cyan) visibility polygon morphing (orange) to another visibility polygon (magenta) $10~{}{\rm m}$ down the future path of the POI (red dotted line). The points for the visibility polygons are the Xs in the corresponding colors. A 2D representation of the urban environment sliced at $5~{}{\rm m}$ above the ground is in blue. The altitude of the visibility polygons is $50~{}{\rm m}$.
@@ 6 Feedback Controller
This section describes a steering controller that drives the Dubins vehicle, system (1), onto the orbits created in Sec. 5.3. First, Sec. 6.1 describes the construction of a constant magnitude vector field using a framework adapted from [54]. The vector field serves to define a reference heading guidance command for the vehicle. Next, Sec. 6.2 derives an expression for the corresponding heading-rate along the vector field. Lastly, Sec. 6.3 proposes a steering controller $u_{\psi}$ with three terms: (i) a proportional term that forces the UAV’s heading to the desired heading based on the vector field, (ii) a feed-forward term that uses the derived heading-rate, and (iii) a term to ensure that the error between the orbit and the UAV approaches zero.
@@ 6.1 Vector Field Guidance for the Desired Heading
The authors in [54] create a generic framework for finding a vector field to attract a single integrator model of a system with state $\mathbb{R}^{n}$ to a time-varying closed trajectory. This work uses that framework to find a specific realization of that vector field for a constant speed single integrator model in $\mathbb{R}^{2}$ to follow the orbit $O$. The resulting vector field generates guidance commands for a steering controller $u_{\psi}$. Consider the single integrator model
@@
where $\Phi(\bm{\xi},t)\in\mathbb{R}$ is a strictly positive function for all values of ${\bm{\xi}}\notin O$ excluding ${\bm{\xi}}\in O$, $V_{1}(a_{1},\ldots,a_{n-1})\in\mathbb{R}^{+}$ is a strictly positive function that depends on the $a_{i}$ described earlier, $H(\bm{\xi},t)\in\mathbb{R}^{+}\backslash\{0\}$ or $H(\bm{\xi},t)\in\mathbb{R}^{-}\backslash\{0\}$ is either a strictly positive or a strictly negative function defining counter-clockwise or clockwise traversal of the trajectory, $\wedge_{i=0}^{n-1}\nabla a_{i}$ is the wedge product [60] between the gradient of the functions $a_{i}$ creating $n$-1-blades, $\prescript{\star}{}{(\cdot)}$ is the Hodge star [60] which maps $n$-1-blades to 1-blades or vectors, $\bm{M}$ is an invertible $n\times n$ matrix where the first $n-1$ rows are $\nabla a_{i}^{\rm T}$ and the last row is $\prescript{\star}{}{\left(\wedge_{i=1}^{n-1}\nabla a_{i}\right)}^{\rm T}$, and the vector $\dot{\bm{\sigma}}$ is $\partial a_{i}/\partial t$ for the first $n-1$ rows and 00 for the last row. In the controller (60), the $\Phi\nabla V$ term determines the direction to the orbit $O$. The controller (60) differs from the notation used in [54]; We have elected to use the Hodge star operator to conform to our convention of using column vectors, as opposed to row vectors, in the wedge product. The next term, $H\prescript{\star}{}{\left(\wedge_{i=1}^{n-1}\nabla a_{i}\right)}$, creates a vector orthogonal to $\Phi\nabla V$ responsible for circulation around the orbit $O$. The final term $\bm{M}^{-1}\dot{\bm{\sigma}}$ is a feed-forward term ensuring the UAV remains on the time-varying orbit $O$. To realize a controller from the framework in [54], this work chooses the function
@@
where $\beta$ is a non-zero positive scalar the controls attracting towards the orbit, and $a_{1}$ is zero on the orbit $O$ for all ${\bm{\xi}}\in O$, and $g_{x},g_{y}$ and $R$ are time-varying terms as introduced earlier. The partial time derivative of the function (61) is
$\frac{\partial a_{1}}{\partial t}=\frac{\beta}{1+\beta^{2}\left(||\bm{\xi}-\bm{g}||-R\right)^{2}}\left(-\dot{{R}}-\frac{\dot{g}_{x}(\xi_{x}-g_{x})+\dot{g}_{y}(\xi_{y}-g_{y})}{||\bm{\xi}-\bm{g}||}\right)$ (62)
The potential function
$V_{1}=\frac{1}{2}a_{1}^{2}$
creates a potential bowl with the bottom of the bowl along the orbit. The feed-forward term $\bm{M}^{-1}\dot{\bm{\sigma}}$ is created by finding the $\bm{M}$ matrix, which involves the gradient
@@
on the first row, where $r=||\bm{\xi}-\bm{g}||$ is the distance between the POI $\bm{g}$ and the inertial position $\bm{\xi}$, and the circulation term is on the second (i.e., last) row,
@@
where the wedge product for one vector is an identity operation $\wedge^{1}_{i=1}\nabla a_{i}=\nabla a_{1}$, and the Hodge star in $\mathbb{R}^{2}$ transforms the directions $\prescript{\star}{}{\bm{e}_{x}}=\bm{e}_{y}$ and $\prescript{\star}{}{\bm{e}_{y}}=-\bm{e}_{x}$. The resulting matrix $\bm{M}$ is
@@
is satisfied for all $\bm{\xi}\in\mathbb{R}^{2}$. When (70) is multiplied by the gradient (66),
@@
To ensure that the magnitude of the controller (60)—composed of (65), (71), (72), and (73)—is equal to the the speed of system (1), solve for the function $H$:
$v^{2}$ $=||\bm{u}||^{2}$
$v^{2}$ $=\left|\left|-\Phi^{\prime}\bm{e}_{r}+\left(\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r}\right)\bm{e}_{r}+H\frac{\beta}{\left(1+\beta^{2}(r-R)^{2}\right)}\bm{e}_{\theta}\right|\right|^{2}$
$v^{2}$ $=\left(-\Phi^{\prime}+\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r}\right)^{2}+H^{2}\left(\frac{\beta}{\left(1+\beta^{2}(r-R)^{2}\right)}\right)^{2}$

### v2

@@ 4.1 Feasibility of Reference Orbits
which is the value of $\kappa$ obtained by substituting (13) into (31), using the identities $\bm{g}\cdot\bm{e}_{r}=v_{g}\cos a$ and $\bm{g}\cdot\bm{e}_{\theta}=v_{g}\sin a$, defining $a=\theta-\gamma$ as the difference between the polar angle and the POI’s direction, introducing the speed ratio $b=\frac{v_{g}}{v}$, defining the ratio between the orbits radial rate and the UAV’s speed $c=\frac{\dot{R}}{v}$, the initial radius $d=R_{0}$, and $R=d+vct$ as the orbit radius at any time $t$. The bounds of the maximization are the bounds on the rate of change of the orbit’s radius from (52), $1\geq b+c$ and $-1\leq-b+c$. The Jacobian of (54) was computed symbolically but is not shown due to space constraints. Figure 3 illustrates the cost function (54) for CCW motion with parameters $d=1$, and $t=0$ in the dark red to white color scale. The curves where elements of the Jacobian are zero also overlayed in Fig. 3 as the blue dashed and green lines.
Figure 3: A plot of the cost function (54) and its Jacobian where $d=1$, $v=1$ and $t=0$ and for varying values of $b$.
In Fig. 3, two local extrema are observed to always occur at $a=\pm\frac{\pi}{2}$ and $c=0$ shown as cyan dots. In Fig. 3, the gray regions are invalid according to Lemma 5. For any set of $d,t$, there is an equivalent value of $d^{\prime},t^{\prime}=0$ such that $cvt+d=cvt^{\prime}+d^{\prime}$; Therefore, evaluating the Jacobian at $t^{\prime}=0$ and $d^{\prime}$ is valid for the corresponding $t>0$ and $d$. Evaluating the Jacobian at $a=\pm\frac{\pi}{2}$ and $c=0$ symbolically results in the zero vector and shows that the critical point is independent of $b$, $d$, and $v$. Thus, the function (54) has local extrema at [±$[\pm\frac{\pi}{2},0]^{\rm T}$ with a value of,
@@
Based on an analysis of plots, the maximum for CCW motion is at $\frac{\pi}{2}$ and $-\frac{\pi}{2}$ for CW motion. ∎
@@
Since $v_{g}\leq v$, then from Lemma 6, the maximum of the right-hand side of (53) is $\kappa=\frac{4}{R}$. This indicates that a UAV with speed $v$ can track a moving POI with any speed $v_{g}\leq v$ as long as the orbit radius is four times greater than the minimum radius $R\geq 4r_{\rm min}$. This is useful because it creates a UAV turn radius selection criterion for a mission.
@@ 4.2 Representing Moving View Volumes via an Adaptive Discretization Approach
VVs for a moving POI remain relatively constant in shape over a road network when the surrounding building geometry also remains relatively unchanged (e.g., when traversing between two spatially uniform structures). However, when one structure ends or changes shape, the VV geometry varies. An adaptive discretization approach chooses an appropriate spacing over the road network for which to compute VVs, such that more VVs are located in regions of greater environmental variability. The adaptive discretization approach relies on quantifying the difference between neighboring VVs $\hat{\alpha}(\bm{g}_{i})$ and $\hat{\alpha}(\bm{g}_{j})$, computed at two points $\bm{g}_{i}$ and $\bm{g}_{j}$, using the difference metric:
@@
The proposed adaptive approach is inspired by a bisection method. The approach increases the number of VVs along a road network edge where the VVs are quickly changing by placing a new VV halfway between two VVs. The approach requires the start and end points of an edge on the road network, $\bm{g}_{i}$ and $\bm{g}_{j}$, a set of initial points along the road network, and a constant threshold parameter $d_{\rm cutoff}$ that is repeatedly compared to (56). The output of the approach is a set of points $\mathcal{P}$ at which the VVs were computed. The approach recursively calls itself twice if the metric (56) is greater than $d_{\rm cutoff}$. The first call updates the endpoint parameter to the middle point, and the second call updates the start point parameter to the middle point (creating two new road segments half the size of the original).
Figure 4: A plot of the VV similarity metric along a path through an urban environment for fixed distance discretization methods ($50~{}{\rm m}$ blue, $25~{}{\rm m}$ orange, $10~{}{\rm m}$ green, and $1~{}{\rm m}$ red) and the adaptive discretization approach (purple) with a cutoff at $d_{\rm cutoff}=500,000$ m3.
Figure 4 plots the distance (56) computed along a path through a city to compare the adaptive discretization with several fixed-spacing discretizations with uniform spacings of $1~{}{\rm m}$, $10~{}{\rm m}$, $25~{}\rm{m}$, $50~{}{\rm m}$. The metric (56) varies greatly for the larger spacings, but the adaptive discretization method maintains the metric at an approximately constant level or below $d_{\rm cutoff}=500,000$ m3 (about 25% of the volume of a hemispherical VV with a radius of 100 m). The value for the adaptive discretization is similar to the $10~{}{\rm m}$ spacing but reduces the spike around $1300~{}{\rm m}$ along the path. A uniform discretization with $1~{}{\rm m}$ spacing creates a refined representation of the VVs, but it requires more computation.
@@ 4.3 Locus of Circular Orbits
Once the path of the POI on the road network has been discretized into a series of $m$ points $\mathcal{P}=\{{\bm{g}}_{0},\ldots,{\bm{g}}_{m-1}\}$ with corresponding VVs $\hat{\alpha}({\bm{g}}_{i})$ and times $t_{0:m-1}=\{t_{0},\ldots,t_{m-1}\}$, then the locus of orbits is created with radii $R_{0:m-1}=\{R_{0},\ldots,R_{m-1}\}$ with centers corresponding to points in $\mathcal{P}$. These radii define the piecewise-linear time-varying radius $R$ as described in Sec. 4.1. The objective (6) is to maximize the standoff distance to the target. Therefore, the radii $R_{i}$ are maximized within the VPs (at altitude $h_{\rm UAV}$) $\hat{\alpha}({\bm{g}}_{i})$ while remaining feasible for the Dubins vehicle.
@@
2: input: $\bm{g}(t)$ the POI trajectory, a set $\mathcal{P}$ of $m$ points along the trajectory $\bm{g}(t)$, $t_{0:m-1}$ the set of times corresponding to the points, $\kappa_{\rm max}$ the maximum curvature of the UAV, $v$ the UAV’s speed
3: output: an ordered set of radii $R_{0:m-1}$ where all points on the orbit can view the corresponding point in $\mathcal{P}$
@@
The algorithm first uses the bisection method MaxRadius to find the largest circle centered on $\bm{g}_{i}$ that can fit inside each VV $\hat{\alpha}({\bm{g}}_{i})$ at height $h_{\rm UAV}$ (Alg. 1, Line 4). The process of finding the largest VO is similar to the maximum inscribed circle problem [25]. The bisection method starts with the minimum radius zero and the maximum sensing distance $d_{\rm max}$ and iteratively increases the minimum and decreases the maximum radius until they are arbitrarily close. The results is a circle that touches the VP at least once. Then, to ensure that the UAV can fly along the morphing orbit, the rate of change of the orbits is checked with to Lemma 5. The first for loop (Alg. 1, Lines 510) iterates forward through the orbits, enforcing the upper bound from (50). The second for loop (Alg. 1, Lines 1116) iterates backwards, enforcing the lower bound from (50). Then, the resulting morphing orbit is checked for curvature constant violations (Alg. 1, Line 18) with Lemma 6. If the check for the curvature of the orbit fails, the path $\bm{g}$ is deemed infeasible for system (1)–(2) with speed $v$ and curvature bound $\kappa_{\rm max}$. An example of the resulting circular orbits can be seen in Fig. 5.
Figure 5: A circular orbit inside the (cyan) visibility polygon morphing (orange) to another visibility polygon (magenta) $10~{}{\rm m}$ down the future path of the POI (red dotted line).
Where the orbit is morphing from the $i$th magenta orbit inside the $i$th VV to the $j$th cyan orbit inside the $j$th VV. The $i$th and $j$th VVs are calculated along the POI’s path, Xs of the corresponding color. The blue footprints of the buildings are shown in Fig 5 and the VPs are calculated at $h_{\rm UAV}=50$ m. Figure 5 illustrates that circular orbits restrict the UAV to a small percentage of the visibility polygon for some environments. Future work may consider more general visibility orbit geometries.
@@ 5 Feedback Controller
This section describes a steering controller that drives the Dubins vehicle, system (1), onto the orbits created in Sec. 4.3. First, Sec. 5.1 describes the construction of a constant magnitude vector field using a framework from [26]. The vector field defines a desired heading guidance command for the vehicle. Next, Sec. 5.2 derives an expression for the heading-rate along the vector field. Lastly, Sec. 5.3 proposes a steering controller $u_{\psi}$ with three terms: (i) a proportional term that forces the UAV’s heading to the desired guidance heading, (ii) a feed-forward term that uses the derived heading-rate, and (iii) a term to ensure that the error between the orbit and the UAV approaches zero.
@@ 5.1 Vector Field Guidance for the Desired Heading
The authors in [26] created a generic framework for finding a vector field to attract a single integrator model of a system with state $\mathbb{R}^{n}$ to a time-varying closed trajectory. This work uses that framework to find a specific realization of that vector field for a constant speed single integrator model in $\mathbb{R}^{2}$ to follow the orbit $O$. The resulting vector field generates guidance commands for a steering controller $u_{\psi}$. Consider the single integrator model
@@
where $\Phi(\bm{\xi},t)\in\mathbb{R}$ is a strictly positive function for ${\bm{\xi}}\notin O$, $V_{1}(a_{1},\ldots,a_{n-1})\in\mathbb{R}^{+}$ is a strictly positive function that depends on the $a_{i}$ described earlier, $H(\bm{\xi},t)\in\mathbb{R}^{+}\backslash\{0\}$ or $H(\bm{\xi},t)\in\mathbb{R}^{-}\backslash\{0\}$ is either a strictly positive or a strictly negative function defining CCW or CW traversal of the trajectory, $\wedge_{i=0}^{n-1}\nabla a_{i}$ is the wedge product [27] between the gradient of the functions $a_{i}$ creating $n$-1-blades, $\prescript{\star}{}{(\cdot)}$ is the Hodge star [27] which maps $n$-1-blades to 1-blades or vectors, $\bm{M}$ is an invertible $n\times n$ matrix where the first $n-1$ rows are $\nabla a_{i}^{\rm T}$ and the last row is $\prescript{\star}{}{\left(\wedge_{i=1}^{n-1}\nabla a_{i}\right)}^{\rm T}$, and the vector $\dot{\bm{\sigma}}$ is $\partial a_{i}/\partial t$ for the first $n-1$ rows and 00 for the last row. In the controller (58), the $\Phi\nabla V$ term is the direction to the orbit $O$. The controller (58) differs from the notation used in [26]; We have elected to use the Hodge star operator to conform to our convention of using column vectors, as opposed to row vectors, in the wedge product. The next term, $H\prescript{\star}{}{\left(\wedge_{i=1}^{n-1}\nabla a_{i}\right)}$, creates a vector orthogonal to $\Phi\nabla V$ for circulation on the orbit $O$. The final term $\bm{M}^{-1}\dot{\bm{\sigma}}$ is a feed-forward term ensuring the UAV remains on the time-varying orbit $O$. To implement a controller from the framework in [26], consider the function
@@
where $\beta$ is a non-zero positive scalar controlling attraction to the orbit, and $a_{1}$ is zero on the orbit $O$ for all ${\bm{\xi}}\in O$, and $g_{x},g_{y}$ and $R$ are time-varying terms as introduced earlier. The partial time derivative of the function (59) is
$\frac{\partial a_{1}}{\partial t}=\frac{\beta}{1+\beta^{2}\left(||\bm{\xi}-\bm{g}||-R\right)^{2}}\left(-\dot{{R}}-\frac{\dot{g}_{x}(\xi_{x}-g_{x})+\dot{g}_{y}(\xi_{y}-g_{y})}{||\bm{\xi}-\bm{g}||}\right)\;.$ (60)
The potential function $V_{1}=\frac{1}{2}a_{1}^{2}$ creates a potential bowl with the bottom of the bowl along the orbit. The feed-forward term $\bm{M}^{-1}\dot{\bm{\sigma}}$ is created by finding the $\bm{M}$ matrix, which involves the gradient
@@
on the first row, where $r=||\bm{\xi}-\bm{g}||$ is the distance between the POI $\bm{g}$ and the inertial position $\bm{\xi}$, and the circulation term
@@
on the second (i.e., last) row where the wedge product for one vector is an identity operation $\wedge^{1}_{i=1}\nabla a_{i}=\nabla a_{1}$, and the Hodge star in $\mathbb{R}^{2}$ transforms the directions $\prescript{\star}{}{\bm{e}_{x}}=\bm{e}_{y}$ and $\prescript{\star}{}{\bm{e}_{y}}=-\bm{e}_{x}$. The resulting matrix $\bm{M}$ is
@@
and is satisfied for all $\bm{\xi}\in\mathbb{R}^{2}$. When (68) is multiplied by the gradient (64),
@@
Solve for $H$ so that the magnitude of the controller (58)—made of (63), (69), (70), and (71)—is the speed of system (1)
$v^{2}=||\bm{u}||^{2}$ $=\left|\left|-\Phi^{\prime}\bm{e}_{r}+\left(\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r}\right)\bm{e}_{r}+H\frac{\beta}{\left(1+\beta^{2}(r-R)^{2}\right)}\bm{e}_{\theta}\right|\right|^{2}$